I need help with this question
A pop growing at an annual rate "r" will triple in a time "t", and is given by the following formula. If the growth rate remains constant and equals 14% per year, how long will it take the population of the time to triple?

Answer by scott8148(6628)

3P=P(1+r)^t __ 3P=P(1.14)^t

dividing by P __ 3=(1.14)^t

taking log __ log(3)=t(log(1.14))

dividing by log(1.14) __ (log(3))/(log(1.14))=t